# On Hamiltonian formulations of the Dirac system

### v1

@@ Abstract
We extend a previously successful discussion of the constrained Schrödinger system through the Dirac–Bergmann algorithm to the case of the Dirac field. In order to follow the analogy, first we discuss the classical Dirac field as a spinorial variable, by introducing properly defined momenta and a suitably modified, factor ordered Poisson bracket. According to the Dirac–Bergmann algorithm two second class Hamiltonian constraints emerge, leading to a factor ordered Dirac bracket on the full phase space. This becomes the Poisson bracket on the reduced phase space in the canonical chart adapted to the shell. The Dirac equation is recovered both as consistency condition on the full phase space and as canonical equation on the reduced phase space. Alternatively, considering the Dirac field as odd Grassmann variable, we present the details of the Dirac–Bergmann algorithm (with either left and righ derivatives acting on Grassmann valued superfunctions and involving a different type of generalized Poisson and Dirac brackets). We propose a recipe for the canonical second quantization of all three versions of the generalized Dirac brackets, yielding the correct fundamental anticommutator.
@@ I Introduction
Most of the physical theories of interest contain constraints, for which the transition from the Lagrangian to the Hamiltonian description requires the Dirac–Bergmann algorithm, developed initially by Dirac DiracConstr and Bergmann with his collaborators (for a historical account see Salisbury and for further details SM ; Sundermayer ; HT ). The full machinery of the formalism through the analysis of a finite-dimensional system has been recently presented in Ref. Brown .
While the corresponding discussions of the Maxwell electrodynamics and of general relativity Wald ; ADM are well known, the Schrödinger field as a constrained system was analyzed relatively late Schrod . There, the Schrödinger equation emerged both as a consistency condition in the full phase space and as canonical equation in the reduced phase space. By implementing the quantization scheme for systems with second class constraints, inconsistencies of previous treatments were eliminated.
@@
For comparing our approach with the interpretation of the classical Dirac field as an odd Grassmann variable, in Section 4 we also carry out the Dirac–Bergmann procedure with both left and righ derivatives acting on Grassmann valued superfunctions. These involve another type of generalized Poisson and Dirac brackets, which are symmetric, rather than antisymmetric, together with definitions of the momenta different from the literature. As far as we are aware of, this discussion was not presented in detail elsewhere.
@@ II.4 Hermitian Lagrangian
$\mathfrak{L}_{IZ}=\frac{i\hbar c}{2}\left(\bar{\psi}\not{\partial}\psi-\not{\partial}\bar{\psi}\psi\right)-mc^{2}\bar{\psi}\psi$ (26)
@@ III The Dirac–Bergmann algorithm
Similarly to the Schrödinger field, either of the Lagrangians (16) or (26) characterizing the dynamics of the Dirac spinor contain linearly the generalized velocities ∂̸$\not{\partial}_{0}\psi∂$ and ∂̸$\not{\partial}_{0}\bar{\psi}∂$, a simptom of a contrained system. Hence the Dirac–Bergmann algorithm needs to be applied for an unambiguous transition from the Lagrangian to the Hamiltonian description.
@@ III.1 Primary Hamiltonian
with a combination of the primary contraints, in which the spinor ∂̸$\not{\partial}_{0}\psi∂$ and the Dirac adjoint spinor ∂̸$\not{\partial}_{0}\bar{\psi}∂$ are unknown functions of the phase space variables.
@@ III.2 Factor ordered Poisson bracket
We introduce a factor ordered Poisson bracket (defined by the requirement that in each term a Dirac adjungate spinor multiplies a spinor) of two phase space functions $f~{}$and $g$ (thus depending only of inner products of Dirac adjungate spinors with spinors) as
@@ III.6 Dirac brackets of canonical data
$\left\{\psi\left(t,\mathbf{r}\right),\bar{\psi}\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=-\frac{i}{\hbar c}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{}.$ (52)
$\left\{\pi\left(t,\mathbf{r}\right),\bar{\pi}\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\frac{i\hbar c}{4}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{}.$ (53)
$\left\{\psi\left(t,\mathbf{r}\right),\psi\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\left\{\bar{\psi}\left(t,\mathbf{r}\right),\bar{\psi}\left(t,\mathbf{r}^{\prime}\right)\right\}=0~{}.$ (54)
@@ III.7 Reduced phase space
$-\frac{1}{2}\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\pi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\psi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\right)$ (62)
@@ IV The Dirac field as Grassmann variable
In an alternative approach explored in Refs. HT and Grassmann2 , classically the Dirac field is regarded as an odd Grassmann variable. For comparison we explicitly present the Dirac–Bergmann algorithm for the Dirac field, treated as an odd Grassmann variable, in two equivalent formulations. For simplifying the notation, we also drop the arguments of the variables in this section.
@@ IV.1 Grassmann variables
In the above notation we have supressed the exterior or wedge product between them. Such products applied repeatedly on the generators lead to superfunctions. By definition, a subclass of superfunctions anticommuting with each other are odd (of parity $1$). Another subclass, consisting of superfunctions commuting with each other and also with the odd ones are even (of parity 00). The simplest superfunctions are the odd (even) Grassmann variables, also dubbed Grassmann numbers, defined as series of odd (even) number products of generators with complex coefficients. Classically the Dirac spinor, its adjoint and Dirac adjoint all qualify as odd Grassman variables, which anticommute, for example $\bar{\psi}\psi=-\psi\bar{\psi}$. By contrast, an even Grassmann variable $E$ commutes with both even or odd Grassman variables, or in fact with any superfunction $F$ as $EF=FE$. Following this approach there is no need for factor ordering, as $\psi,\ \psi^{{\dagger}},\ \bar{\psi}$ appear on equal footing (they can be thought of as $4\times 4$ matrices). The price to pay is that when calculating the brackets of the superfunctions, two versions of partial derivatives with respect to the generators need to be introduced. These are the left (denoted $L$, acting from the left to the right) and right ($R$, acting from the right to the left) derivatives, related as
@@ IV.2 Hamiltonian formalism in terms of left derivatives
They differ by $\gamma^{0}$ factors in the denominator from the definitions advanced in Ref. HT (which contain $\partial_{0}$ instead ∂̸$\not{\partial}_{0}∂$). The purpose of this change is to remain as close as possible to the expressions of the momenta introduced in the previous chapter, Eqs. (27). Indeed, $\bar{\pi}_{L}$ has the same form as $\bar{\pi}$, however $\pi_{L}$ differs in a minus sign. The momenta are odd variables.
@@
After accounting for the change in the definitions of momenta, these agree with the corresponding equations presented in Ref. HT . Note that the left derivatives above can also be rewritten as right derivatives through Eq. (71).
@@ IV.4 Dirac-Bergmann algorithm
$\{F,G\}_{D+}^{L,R}=\{F,G\}_{+}^{L,R}-\int\mathrm{d}\mathbf{r}^{\prime\prime}\mathrm{d}\mathbf{r}^{\prime\prime\prime}\{f,\phi^{i}\}_{+}^{L,R}(A_{ij}^{L,R})^{-1}\{\phi^{j},g\}_{+}^{L,R}~{}.$ (94)
@@
Note that now only the Dirac brackets of the canonical pairs are identical when calculated in terms of left or right derivatives. Hence, the $L,R$ superscripts may be omitted from both the Poisson and Dirac brackets of the canonical pairs only. Also note the appearance of the $1/2$ factors, similarly to the Dirac brackets of the canonical pairs in the spinor formalism.
@@ IV.5 Reduced phase space of Grassmann variables
on the reduced phase space and the pair of contraints
@@ V.3 Grassmannian approach with right derivatives
VI Concluding Remarks
@@
We generalized a previously successful discussion of the Schrödinger field Schrod as a constrained system through the Dirac–Bergmann algorithm to the case of the Dirac field. This was motivated by the analogy of the Lagrangians describing the Dirac field with the simplest Schiff and Hermitian Henley–Thirring Lagrangians of the Scrödinger field.
@@
In order to follow as closely as possible the treatment of the Schrödinger field, we discussed the Dirac field classically as a spinor. Then, for the Hamiltonian evolution we introduced suitably modified, factor ordered Poisson and Dirac brackets (in which Dirac adjungate spinors are always followed by spinors) and we recovered all essential ingredients, similarly to the case of the Schrödinger field. We introduced momenta, which are Dirac adjoints of the respective spinorial field variables. This ensured that no $\gamma^{0}$ factors appear in the Poisson bracket.
@@
When the Dirac field is considered an odd Grassmann variable, it obeys the same algebraic structure classically as in the quantum version of the theory. Hence, we carried out the Dirac–Bergmann algorithm for the Grassmannian Dirac field with both left and righ derivatives acting on Grassmann valued superfunctions. These involved other types of generalized Poisson and Dirac brackets, symmetric in the canonical data, rather than antisymmetric.

### v2

@@ Abstract
We extend a previously successful discussion of the constrained Schrödinger system through the Dirac–Bergmann algorithm to the case of the Dirac field. In order to follow the analogy, first we discuss the classical Dirac field as a spinorial variable, by introducing properly defined momenta and a suitably modified, factor ordered Poisson bracket. According to the Dirac–Bergmann algorithm two second class Hamiltonian constraints emerge, leading to a factor ordered Dirac bracket on the full phase space. This becomes the Poisson bracket on the reduced phase space in the canonical chart adapted to the shell. The Dirac equation is recovered both as consistency condition on the full phase space and as canonical equation on the reduced phase space. Alternatively, considering the Dirac field as odd Grassmann variable, we present the details of the Dirac–Bergmann algorithm (with either left and right derivatives acting on Grassmann valued superfunctions and involving a different type of generalized Poisson and Dirac brackets). We propose a recipe for the canonical second quantization of all three versions of the generalized Dirac brackets, yielding the correct fundamental anticommutator.
@@ I Introduction
Most of the physical theories of interest contain constraints, for which the transition from the Lagrangian to the Hamiltonian description requires the Dirac–Bergmann algorithm, developed initially by Dirac DiracConstr and Bergmann with his collaborators (for a historical account see Salisbury and for further details SM ; Sundermeyer ; HT ). The full machinery of the formalism through the analysis of a finite-dimensional system has been recently presented in Ref. Brown .
Geometric generalizations of the Dirac–Bergmann algorithm have also been developed. The algorithm by Gotay, Nester, and Hinds in Ref. GNH78 employs a global presymplectic structure, generalizing and improving on the Dirac–Bergmann algorithm, as the latter can be regarded as its local version. In the companion paper GN79 , Gotay and Nester further emphasize that while the Dirac–Bergmann results cannot always be transferred to the velocity phase space, their method based on the presymplectic structure allows that. This was later investigated for a broad class of admissible Lagrangians, defined as those which allow for second-order Euler–Lagrange equations on certain submanifolds of the velocity phase space GN80 . A multisymplectic approach was also developed in Ref. multisymp and explored in the discussion of various constrained systems, including Einstein–Hilbert multisympEH and metric-affine multisympEP gravity theories, also more recently for scalar fields, Chern-Simons gravity, and bosonic string theory multisympField .
While the Dirac–Bergmann algorithm applied for Maxwell electrodynamics or general relativity Wald ; ADM are well known, the Schrödinger field as a constrained system was analysed relatively late Schrod . There, the Schrödinger equation emerged both as a consistency condition in the full phase space and as canonical equation in the reduced phase space. By implementing the quantization scheme for systems with second class constraints, inconsistencies of previous treatments were eliminated.
@@
For comparing our approach with the interpretation of the classical Dirac field as an odd Grassmann variable, in Section 4 we also carry out the Dirac–Bergmann procedure with both left and right derivatives acting on Grassmann valued superfunctions. These involve another type of generalized Poisson and Dirac brackets, which are symmetric, rather than antisymmetric, together with definitions of the momenta different from the literature. As far as we are aware of, this discussion was not presented in detail elsewhere.
@@
Section 6 contains a brief comparison of our results with the flat limit of previous discussions of fermionic (Dirac) systems in gravitational field. Then we present our concluding remarks in Section 7.
@@ II.4 Hermitian Lagrangian
$\mathfrak{L}_{IZ}=\frac{i\hbar c}{2}\left(\bar{\psi}\not{\partial}\psi-\not{\partial}\bar{\psi}\psi\right)-mc^{2}\bar{\psi}\psi~{}.$ (26)
@@ III The Dirac–Bergmann algorithm
Similarly to the Schrödinger field, either of the Lagrangians (16) or (26) characterizing the dynamics of the Dirac spinor contain linearly the generalized velocities ∂̸$\not{\partial}_{0}\psi∂$ and ∂̸$\not{\partial}_{0}\bar{\psi}∂$, a simptom of a constrained system. Hence the Dirac–Bergmann algorithm needs to be applied for an unambiguous transition from the Lagrangian to the Hamiltonian description.
@@ III.1 Primary Hamiltonian
with a combination of the primary constraints, in which the spinor ∂̸$\not{\partial}_{0}\psi∂$ and the Dirac adjoint spinor ∂̸$\not{\partial}_{0}\bar{\psi}∂$ are unknown functions of the phase space variables.
@@ III.2 Factor ordered Poisson bracket
We introduce a factor ordered Poisson bracket (defined by the requirement that in each term a Dirac adjoint spinor multiplies a spinor) of two phase space functions $f~{}$and $g$ (thus depending only of inner products of Dirac adjoint spinors with spinors) as
@@ III.6 Dirac brackets of canonical data
$\left\{\psi\left(t,\mathbf{r}\right),\bar{\psi}\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=-\frac{i}{\hbar c}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{},$ (52)
$\left\{\pi\left(t,\mathbf{r}\right),\bar{\pi}\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\frac{i\hbar c}{4}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{},$ (53)
$\left\{\psi\left(t,\mathbf{r}\right),\psi\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\left\{\bar{\psi}\left(t,\mathbf{r}\right),\bar{\psi}\left(t,\mathbf{r}^{\prime}\right)\right\}=0~{},$ (54)
@@ III.7 Reduced phase space
$-\frac{1}{2}\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\pi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\psi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\right)~{}.$ (62)
@@ IV The Dirac field as Grassmann variable
In an alternative approach explored in Refs. HT and Grassmann2 , classically fermionic fields are regarded as odd Grassmann variables, nevertheless these references do not present the case of the Dirac field explicitly. For comparing with our earlier treatment, we explicitly present the Dirac–Bergmann algorithm for the Dirac field, treated as an odd Grassmann variable, in two equivalent formulations. For simplifying the notation, we also drop the arguments of the variables in this section.
@@ IV.1 Grassmann variables
In the above notation we have suppressed the exterior or wedge product between them. Such products applied repeatedly on the generators lead to superfunctions. By definition, a subclass of superfunctions anticommuting with each other are odd (of parity $1$). Another subclass, consisting of superfunctions commuting with each other and also with the odd ones are even (of parity 00). The simplest superfunctions are the odd (even) Grassmann variables, also dubbed Grassmann numbers, defined as series of odd (even) number products of generators with complex coefficients. Classically the Dirac spinor, its adjoint and Dirac adjoint all qualify as odd Grassman variables, which anticommute, for example $\bar{\psi}\psi=-\psi\bar{\psi}$. By contrast, an even Grassmann variable $E$ commutes with both even or odd Grassman variables, or in fact with any superfunction $F$ as $EF=FE$. Following this approach there is no need for factor ordering, as $\psi,\ \psi^{{\dagger}},\ \bar{\psi}$ appear on equal footing (they can be thought of as $4\times 4$ matrices). The price to pay is that when calculating the brackets of the superfunctions, two versions of partial derivatives with respect to the generators need to be introduced. These are the left (denoted $L$, acting from the left to the right) and right ($R$, acting from the right to the left) derivatives, related as
@@ IV.2 Hamiltonian formalism in terms of left derivatives
They differ by $\gamma^{0}$ factors in the denominator from the expressions obtainable by applying the definitions advanced in Ref. HT (which contain $\partial_{0}$ instead ∂̸$\not{\partial}_{0}∂$). The purpose of this change is to remain as close as possible to the expressions of the momenta introduced in the previous chapter, Eqs. (27). Indeed, $\bar{\pi}_{L}$ has the same form as $\bar{\pi}$, however $\pi_{L}$ differs in a minus sign. The momenta are odd variables.
@@
After accounting for the change in the definitions of momenta, these agree with the expressions obtainable through the canonical equations given in Ref. HT . Note that the left derivatives above can also be rewritten as right derivatives through Eq. (71).
@@ IV.4 Dirac-Bergmann algorithm
$\{F,G\}_{D+}^{L,R}=\{F,G\}_{+}^{L,R}-\int\int\mathrm{d}\mathbf{r}^{\prime\prime}\mathrm{d}\mathbf{r}^{\prime\prime\prime}\{f,\phi^{i}\}_{+}^{L,R}(A_{ij}^{L,R})^{-1}\{\phi^{j},g\}_{+}^{L,R}~{}.$ (94)
@@
The relations 97 containing left derivatives have also been given in Ref. Sundermeyer , however in different notation (our momenta differ by factors of $\gamma^{0}$). Note that now only the Dirac brackets of the canonical pairs are identical when calculated in terms of left or right derivatives. Hence, the $L,R$ superscripts may be omitted from both the Poisson and Dirac brackets of the canonical pairs only. Also note the appearance of the $1/2$ factors, similarly to the Dirac brackets of the canonical pairs in the spinor formalism.
@@ IV.5 Reduced phase space of Grassmann variables
on the reduced phase space and the pair of constraints
@@ V.2 Grassmannian approach with left derivatives
After the changes in notation, this quantization scheme is equivalent to the one proposed by Casalbuoni Casalbuoni .
@@ V.3 Grassmannian approach with right derivatives
VI On the coupled gravity-Dirac field system
Instead of a metric description, suitable to deal with bosonic (tensorial) fields, gravity can also be described in terms of 16 tetrad variables, as worked out in 1929 by Weyl Weyl29 . The four tetrad vectors (with 4 spacetime components each) represent local orthonormal Lorentz frames at each space-time point with respect to which spinors can be defined. This choice of gravitational variables naturally allows for covariant derivatives of spinors, in which the spin connection (or Ricci rotation coefficients, with 24 components, having two tetrad indices) couples to the $\gamma^{a}$ matrices.
The canonical (Arnowitt—Deser—Misner type) formulation of the dynamics of gravity coupled to the Dirac field has been studied long time ago. As early as 1963 Kibble discussed the canonical form of the Lagrangian describing the interacting system of gravity and Dirac fields, in the framework of the tetrad formalism of gravity Kibble63 . He derived the dynamics of the Dirac field from a sum of the simplest Lagrangian density $\mathfrak{L}_{BD}$ with a spin–spin interaction Lagrangian (bearing similarity to the one arising in the Fermi interaction). Our analysis, by contrast, proceeds from the Hermitian Lagrangian density $\mathfrak{L}_{IZ}$. Kibble assumed anticommuting Dirac spinors, similarly as in our discussion employing anticommuting Grassmann variables. In the first part of our treatment however, we introduced factor ordering instead (requiring Dirac adjoint spinors to multiply Dirac spinors, in this order).
Deser and Isham in Ref. DeserIsham76 provided a crystal-clear canonical description of the gravitational sector in terms of tetrad variables, in which the role of the six pairs of geometrodynamical variables (the induced metric components and associated momenta, the trace-modified extrinsic curvature) is taken by the twelve pairs of variables related to the spatial tetrad components and their respective associated momenta. While the six pairs of geometrodynamical variables can be reduced to two pairs (as the first class Hamiltonian and diffeomorphism constraints each reduce the degrees of freedom by two Sundermeyer ), in the tetrad formalism additional degrees of freedom emerge due to the six-parameter Lorentz transformations of the tetrad. Each of the related additional six angular momentum type constraints, being also first class, reduce the degrees of freedom by two, reestablishing two pairs of canonical variables as physical. The Lorentz invariance of the gravitational action was re-established by inserting the angular momentum type constraints into the action with a set of six Lagrange multipliers. Although the canonical Dirac action for a coupled Dirac field has not been given explicitly, no other obstacle for that was envisioned than the tedious nature of the involved calculations.
Both the Einstein and Dirac Lagrangians, also the respective equations of motion were given in terms of a suitably chosen spinor Lagrangian in tetrads suited to the 3+1 formalism of general relativity GeheniauHenneaux77 , reproducing and generalizing earlier results by Dirac DiracGS . There, the spinor Lagrangian is the generalization of $\mathfrak{L}_{IZ}$ to curved spacetime, however the Hamiltonian given in the Note Added in Proof does not reproduce our result.
In Ref. NelsonTeitelboim78 Nelson and Teitelboim gave the action for interacting gravitational and Dirac fields expressed in canonical form, in tetrad and spinorial variables. The results of Ref. DeserIsham76 were reproduced, supplementing the Hamiltonian and diffeomorphism constraints with the contributions from the Dirac spinor. The emerging Dirac Hamiltonian in the flat limit $L^{i}\mapsto\gamma^{i}$ does not agree with our Hamiltonian (29). Agreement can be reached only at the price of inserting the definitions of the momenta in the Hamiltonian (a forbidden step before variation), and by dropping a spurious factor $i$ from the mass term of the Hamiltonian given in Ref. NelsonTeitelboim78 . The Dirac action in Ref. NelsonTeitelboim78 is again the covariant version of $\mathfrak{L}_{IZ}$ (after omitting a spurious factor $i$ from the mass term), however the authors do not allow the variation with respect to the Dirac adjoint spinor. The discussion then proceeded in terms of the real and imaginary parts of the spinor field, hence the comparison with our methods is not immediate. Solely the brackets of the spinor field and its Dirac adjoint are given, reproducing our Dirac bracket (52) in the flat limit, however the Dirac brackets (57-58) between the canonical variables, which are the most important in our discussion, are missing.
Therefore, our discussion goes beyond the flat spacetime limit of the results of the references mentioned above.
It is well known that gravity can be also described in terms of torsion Hayashi ; Hayashi2 or nonmetricity NewerGR , recently summarized in Ref. Heinsenberg1 . The three approaches are sometimes referred as the geometric trinity of gravity Trinity . It has been proven by Koivisto, that the Dirac fermions described by the covariant version of the Hermitian Lagrangian $\mathcal{L_{IZ}}$ are unaffected by non-metricity Koivisto . Moreover, the consistency of the coupling to the torsion has been disputed Obukhov . Hence, when the connection does not have a contorsion part, but it has a disformation part (related to its nonmetricity $Q_{abc}$), Dirac fermions do not couple to the latter and their minimal coupling to any torsionfree gravity theory is viable HeisenbergJCAP ; telecoupling .
Modified gravity generalizations of both $f(R)$ (see the monography HarkoLobo with references therein, also LeviSaid ) and $f(Q)$ types (see Refs. LazkozQ ; BoehmerQ ; HeisenbergQ ) have been discussed extensively. Dirac fields in $f(Q)$ theories were investigated in Ref. CarloniQ . In principle, $f(R,Q)$ type gravity models could also properly incorporate minimally coupled Dirac spinors.
VII Concluding Remarks
@@
We generalized a previously successful discussion of the Schrödinger field Schrod as a constrained system through the Dirac–Bergmann algorithm to the case of the Dirac field. This was motivated by the analogy of the Lagrangians describing the Dirac field with the simplest Schiff and Hermitian Henley–Thirring Lagrangians of the Schrödinger field.
@@
In order to follow as closely as possible the treatment of the Schrödinger field, we discussed the Dirac field classically as a spinor. Then, for the Hamiltonian evolution we introduced suitably modified, factor ordered Poisson and Dirac brackets (in which Dirac adjoint spinors are always followed by spinors) and we recovered all essential ingredients, similarly to the case of the Schrödinger field. We introduced momenta, which are Dirac adjoints of the respective spinorial field variables. This ensured that no $\gamma^{0}$ factors appear in the Poisson bracket.
@@
When the Dirac field is considered an odd Grassmann variable, it obeys the same algebraic structure classically as in the quantum version of the theory. Hence, we carried out the Dirac–Bergmann algorithm for the Grassmannian Dirac field with both left and right derivatives acting on Grassmann valued superfunctions. These involved other types of generalized Poisson and Dirac brackets, symmetric in the canonical data, rather than antisymmetric.